What are the dimensions of a rectangle with a perimeter of 200 if the length is 4 times width ?
Answer by mananth(16946)

width = x
length = 4x
perimeter = 2*(L+W)
200=2*(5x)
200=10x
/10
x=20
length = 20*4=80
width = 20
length = 80